# Byzantine Reliable Broadcast with Low Communication and Time Complexity

### v1

@@ 7. Related Work
As mentioned in §1, the first reliable broadcast algorithm, achieving a communication complexity of $O(|m|n^{2})$, was presented by Bracha (Bracha, 1987). An algorithm with a much improved bound of $O(|m|n+\kappa n^{2}\log(n))$, based on the use of erasure codes, was published 18 years later by Cachin and Tessaro (Cachin and Tessaro, 2005), making use of collision-resistant hash functions of size $\kappa$. By contrast, Bracha’s algorithm is error-free, i.e., it does not rely on any cryptographic assumptions.
@@
It has also been shown how to achieve a communication complexity of $O(n|m|+\kappa n^{2})$ using only collision-resistant hash functions (Das et al., 2021), improving upon the algorithm by Cachin and Tessaro. The downside of this algorithm is that it is not balanced and it has a higher computational cost. Algorithm $\mathcal{A}_{bit}$ of §3 falls into the category of reliable broadcast algorithms that rely on collision-resistant hash functions as well.
There are reliable broadcast algorithms that require other cryptographic primitives. For example, an algorithm has been proposed that achieves a communication complexity of $O(n|m|+\kappa n^{2})$ but requires a trusted setup for a public key infrastructure and cryptographic accumulators (Nayak et al., 2020). This bound has been improved to $O(n|m|+\kappa n+n^{2})$ using threshold signatures (Alhaddad et al., 2022). The algorithm $\mathcal{A}_{sig}$ defined in §4 uses threshold signatures to achieve an optimal (good-case) time complexity.
The best known upper bounds, with and without cryptographic assumptions, are almost asymptotically tight considering the lower bound of $\Omega(n|m|+n^{2})$ (Dolev and Reischuk, 1985). The algorithms in this paper improve the effective communication complexity for large messages by reducing the constant of the first term. Moreover, most error-free algorithms, excluding Bracha’s algorithm, and also the algorithms that use more cryptographic tooling than hash functions, have a higher time complexity than $\mathcal{A}_{bit}$ and $\mathcal{A}_{sig}$, the latter of which having an optimal time complexity (Abraham et al., 2021b).
@@ 8. Conclusion
The mechanism introduced in this paper lowers the communication complexity of Byzantine reliable broadcast for large messages. The presented algorithms, which utilize this mechanism, further guarantee near-optimal and optimal bounds on the time complexity while also keeping the space complexity close to optimal. As mentioned in §1, numerous applications make use of reliable broadcast as a subroutine. Therefore, it may be worthwhile to revisit selected applications to determine if the techniques introduced in this paper can be used to obtain stronger results.

### v2

@@ 6. Related Work
As mentioned in §1, the first reliable broadcast algorithm, achieving a communication complexity of $O(|m|n^{2})$, was presented by Bracha (Bracha, 1987). An algorithm with a much improved bound of $O(|m|n+\kappa n^{2}\log(n))$, making use of erasure codes, was published 18 years later by Cachin and Tessaro (Cachin and Tessaro, 2005), making use of collision-resistant hash functions of size $\kappa$. By contrast, Bracha’s algorithm is error-free, i.e., it does not rely on any cryptographic assumptions.
@@
It has also been shown how to achieve a communication complexity of $O(n|m|+\kappa n^{2})$ using only collision-resistant hash functions (Das et al., 2021), improving upon the algorithm by Cachin and Tessaro. The downside of this algorithm is that it is not balanced and it has a higher computational cost. Algorithm $\mathcal{A}_{bit}$ introduced in §3 falls into the category of reliable broadcast algorithms that rely on collision-resistant hash functions as well. There are reliable broadcast algorithms that require other cryptographic primitives. For example, an algorithm has been proposed that achieves a communication complexity of $O(n|m|+\kappa n^{2})$ but requires a trusted setup for a public key infrastructure and cryptographic accumulators (Nayak et al., 2020). This bound has been improved to $O(n|m|+\kappa n+n^{2})$ using threshold signatures (Alhaddad et al., 2022). The algorithm $\mathcal{A}_{sig}$ defined in §4 uses threshold signatures to achieve an optimal (good-case) time complexity.
The best known upper bounds, with and without cryptographic assumptions, are almost asymptotically tight considering the lower bound of $\Omega(n|m|+n^{2})$ (Dolev and Reischuk, 1985). The algorithms in this paper improve the effective communication complexity for large messages by reducing the constant of the first term. Moreover, most error-free algorithms, excluding Bracha’s algorithm, and also the algorithms that use more cryptographic tooling than hash functions, have a higher time complexity than $\mathcal{A}_{bit}$ and $\mathcal{A}_{sig}$, the latter having an optimal time complexity (Abraham et al., 2021b).
@@ 7. Conclusion
The mechanism introduced in this paper lowers the communication complexity of Byzantine reliable broadcast for large messages. The presented algorithms, which utilize this mechanism, further guarantee near-optimal and optimal bounds on the time complexity while also keeping the space complexity close to optimal. As mentioned in §1, numerous applications make use of reliable broadcast as a subroutine. Therefore, it may be worthwhile to revisit selected applications to determine if the presented techniques can be used to obtain stronger results.